Out of ten students the favorite drink of 7 is Cokeand the favorite drink of the rest is fanta.one of the students is asked at random what is the probability that the favorite drink of the student is
a .Coke b.fanta c.neither Coke or fanta
d.either Coke or fanta
1
Expert's answer
2021-09-30T15:42:50-0400

a. Total number of outcomes: "n=10" ;

Favorable outcomes: "m=7" .

Then the wanted probability is

"p = \frac{m}{n} = \frac{7}{{10}} = 0.7"

Answer: "p=0.7"

b. Total number of outcomes: "n=10" ;

Favorable outcomes: "m=10-7=3" .

Then the wanted probability is

"p = \frac{m}{n} = \frac{3}{{10}} = 0.3"

Answer: "p=0.3"

c. Since there are no students among the students whose favorite drink is neither coca nor fanta, the wanted probability is "p=0" .

Answer: "p=0"

d. Since every student likes either a cook or a fanta, the wanted probability is "p=1" .

Answer: "p=1"